OpenStudy (anonymous):

I'll give a medal and fan whoever can help me! Solve the equation 2=-4+4 sqrt(x^3) round your answer to the nearest hundredth

OpenStudy (photon336):

\[2 = -4+4\sqrt{x^{3}}\]

OpenStudy (photon336):

this is your equation right>?

OpenStudy (photon336):

Our first objective is to get x by itself \[2+4 = 4\sqrt{x^{3}}\]

OpenStudy (anonymous):

Yeah that's my equation. I think the answer is 1.31

OpenStudy (anonymous):

I already sent it over though so we'll see how it goes.

OpenStudy (photon336):

Just incase you want to know how to do this step by step. \[6 = 4\sqrt{x^{3}}\] \[\frac{ 6 }{ 4 } = \sqrt{x^{3}}\] \[(\frac{ 6 }{ 4 })^{2} = (\sqrt{x^{3}})^{2}\] \[\frac{ 36 }{ 16 } = x^{3}\] \[\sqrt[3]{\frac{ 36 }{ 16 }} = x\]

OpenStudy (anonymous):

Okay, thank you